OpenStudy (anonymous):

6^x=216 How do you solve x ?

OpenStudy (luigi0210):

you could either set it up as a log function and plug it into the calculator or get like bases.

OpenStudy (anonymous):

without a calculator

OpenStudy (anonymous):

Thats easy ,Just move the power to the other side by its inverse Like 6^x=216 , x=216^1/6 and then calculator

OpenStudy (anonymous):

or sixth root 216

OpenStudy (anonymous):

Thanks... x is 3 right ?

OpenStudy (luigi0210):

\[6^x=6^3\] so yes x =3